A 5,5 kg ball is thrown vertically down at 2,5 m/s from height of 2 m above the ground . If the coefficient of restitution between the ground and ball is 0.3 how high shall the ball rise after striking the ground
1
Expert's answer
2021-07-20T09:55:41-0400

Gives

M=5.5kg

V=2.5m/sec

Hight(H) =2m

Restitution cofficient (e) =0.3

"V_i=\sqrt{2gH}=\sqrt{2\times9.8\times2}=\sqrt{39.2}=6.26m/sec"

"V_f=eV_i"


"h=\frac{V_f^2}{2g}=\frac{(eV_i)^2}{2g}=\frac{e^2\times (\sqrt{4g})^2}{2g}=2e^2=2\times0.3^2=0.18m"